Determine the general solution to the exact differential equation

(e2y-ycosxy)dx+(2xe2y-xcosxy+2y)dy=0


1
Expert's answer
2021-08-05T17:16:19-0400
"\dfrac{\partial M}{\partial y}=2e^{2y}-\cos(xy)+xy\sin(xy)"

"\dfrac{\partial N}{\partial x}=2e^{2y}-\cos(xy)+xy\sin(xy)"

"\dfrac{\partial M}{\partial y}=\dfrac{\partial N}{\partial x}"


"\dfrac{\partial u}{\partial x}=e^{2y}-y\cos(xy)"

"\dfrac{\partial u}{\partial y}=2xe^{2y}-x\cos(xy)+2y"

"u(x,y)=\int(e^{2y}-y\cos(xy))dx+\varphi(y)"

"=xe^{2y}-\sin(xy)+\varphi(y)"

"\dfrac{\partial u}{\partial y}=2xe^{2y}-x\cos(xy)+\varphi'(y)"

"=2xe^{2y}-x\cos(xy)+2y"

"\varphi'(y)=2y"

"\varphi(y)=y^2+C_1"

"u=xe^{2y}-\sin(xy)+y^2+C_1"


"xe^{2y}-\sin(xy)+y^2=C"